Yonela purchased a freezer at a cost of R285,000.00. The estimated life of the freezer is 5 years. At which time they have no salvage value?

the company would like to compare allowable depreciation methods and decide to prepare depreciation method schedules for the for the fixtures using the straight-line doble declining-balance and sum-of the year digit method of depreciation schedules.

1. What is the book value at the end of the third year using the striaght line depreciation method.
2. Using the double-declining-balance method of depreciation, what is the book value at the end of the 3rd year?
3. With the sum of year-digit method of depreciation what is the cumulative depreciation at the end of the 3rd year?
4. What amount of depreciation will be taken in four years with each method?
1
Expert's answer
2020-09-10T18:39:44-0400

A company purchased factory equipment for $700,000. It is estimated that the equipment will have a $70,000 salvage value at the end of its estimated 5-year useful life. If the company uses the double-declining-balance method of depreciation, the amount of annual depreciation recorded for the second year after purchase would be

1) ($700,000 - 0) × .40 = $280,000; ($700,000 - $280,000) × .40 = $168,000


2) $85,000 + $3,500 + $10,000 = $98,500; [($98,500 - $15,000) / 5] × 2 = $33,400


3)($400,000 - $25,000) / 15,000 = $25; $3,300 × $25 = $82,500


4)[($324,000 - $24,000) / 5] × 6/12 = $30,000